OpenStudy (anonymous):

Hey lalaly, I have one more: Factor the expression 16a^2-72ab+81b^2 into a product of binomials. Can you help me with this type of equation please?:)

OpenStudy (diyadiya):

(4a-9b)^2 (a-b)^2=a^2-2ab+b^2

OpenStudy (lalaly):

\[a^2-2ab+b^2= (a-b)(a-b)\] \[16a^2-72ab+81b^2= (4a-9b)(4a-9b)\]

OpenStudy (lalaly):

notice that \[\large{16a^2-72ab+81b^2}= (4a)^2-(2 \times 4 \times 9)ab+ (9b)^2\]

OpenStudy (anonymous):

so it would be (4a-9b)^2 = (4a-9b)(4a-9b), am I right?

OpenStudy (lalaly):

yess thats right:D

OpenStudy (anonymous):

oh you got this:) great studying

OpenStudy (lalaly):

;):D

OpenStudy (anonymous):

;-)